2.4g of a compound of coarbon, hydrogen and oxygen gave on combustion 3.52g of co2 and 1.44g of h2o.The Mr of the compound was found to be 60. What are the masses of c H o and find the empirical formula and the molecular of the formula
1
Expert's answer
2020-11-04T14:13:46-0500

1) moles C

"3.52 g (CO_2)\times \frac{1 mole (CO_2)}{44.01 g (CO_2)}\times \frac{1 mole (C)}{1 mole (CO_2)}= 0.0800 moles (C)"

mass C

"0.0800 moles (C)\times \frac{12.01 g (C)}{1 mole (C)}=0.961 g(C)"


2) moles H

"1.44 g (H_2O)\times\frac{1 mole (H_2O)}{18.02 g (H_2O)}\times \frac{2 mole (H)}{1 mole (H_2O)}=0.160 moles (H)"

mass H

"0.160 moles (H)\times \frac{1.01 g(H)}{1 mole (H)}= 0.161 g (H)"


3) mass O

"2.4=m(O)+m(C)+m(H)"

"2.4= m(O) +0.961 + 0.161"

m"(O) = 1.278g"


moles O

"1.278 g (O) \times \frac{1 mole (O)}{16.00 g (O)}=0.08 moles (O)"


4) Empirical formula

"n(C):n(H):n(O) = 0.08:0.16:0.08=1:2:1"

"EF=CH_2O"


5) Molecular formula

M"(CH_2O)=30.03\frac{g}{mol}"

"\frac{M(MF)}{M(EF)}=\frac{60}{30}=2"


"MF=EF\times2= C_2H_4O_2"